OpenStudy (anonymous):

Given that x⊙y=x+4⋅y and that x⊕y=x-4⋅y Evaluate this expression: (5⊙(1⊙(-1)))⊕(-5)

myininaya (myininaya):

lets do the most inside operation first we have x=1 and y=-1 1+4(-1)=1-4=-3 now we have x=5 and y=-3 5-4(-3)=5+12=17 now we have x=17and y=-5 and we are using the last operation 17-4(-5)=17+20=37

myininaya (myininaya):

oops i messed up on the second one we have x=5 and y=-3 5+4(-3)=5-12=-7 now we have x=-7 and y=-5 -7-4(-5)=-7+20=13

OpenStudy (anonymous):

should be 13

myininaya (myininaya):

yes i know i used the wrong operation for the second one

OpenStudy (anonymous):

(5+4*(1+4*(-1)))-4*(-5)=(5+4*(-3))+20=5-12+20=13